If (2x - 1)(2 + 5x)=Ax2 + Bx + C, what is the value of B?
the Ax2 means dat 2 is da power of A;

Answer by stanbon(75887)

If (2x - 1)(2 + 5x)= Ax2 + Bx + C, what is the value of B?
-----------
Find the coefficient of "x"
4x-5x = -1x
B= -1
Cheers,
Stan H.